OpenStudy (anonymous):

An urn contains 10 white balls and 5 blue balls. Draws are made repeatedly from the urn as follows. On each draw, a ball is drawn and its color noted; then it is replaced in the urn along with 3 more balls of its color. a. Find the chance that the first ball drawn is blue. b. Find the chance that the second ball drawn is blue, given that the first ball drawn is blue. c. Find the chance that the second ball drawn is blue. d. Find the chance that the first ball drawn is blue, given that the second ball drawn is blue.

OpenStudy (anonymous):

a. bb/ wb+bb b. bb+increment / wb+bb+increment c. (1st draw bb x 2nd draw bb) + (1st draw wb x 2nd draw bb) increment on both 2nd draws since you know what was color was the 1st draws d. use baye's backward if u know what it is

OpenStudy (anonymous):

a. done (1/3) b. done (4/9) c. i cannot understand the explanation :( d. i know bayes teorem, but im not sure how to apply in this point. a hint please

OpenStudy (amistre64):

for c, determine the number of ways to pull out 3 balls then see how many of them have a blue ball on the second pull

OpenStudy (amistre64):

err, pull out 2 balls that is :)

OpenStudy (anonymous):

Solution please :) c and d :)

OpenStudy (amistre64):

thats not how this site is spose to work. we guide, you try, we adjust

OpenStudy (amistre64):

im pretty sure that since the balls are returned, that the second ball blue is the same chance as a first ball blue .... but i could be wrong on that

OpenStudy (amistre64):

...oh, replaced with 3 more of its color .... thats cheating lol

OpenStudy (amistre64):

since the first pull drops in 3 more of the same color; w, then b; 10/15 * 5/18 b, then b; 5/15 * 8/18 and add em up sounds about right

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

c. 1/3 :)

OpenStudy (anonymous):

and the last one please !!!!

OpenStudy (anonymous):

bayes teorem

OpenStudy (amistre64):

any way to determine if these are correct solutions? just because its been awhile since i tried to think through this stuff

OpenStudy (amistre64):

ww = 10/15 * 13/18 wb = 10/15 * 5/18 bw = 5/15 * 10/18 bb = 5/15 * 8/18 wb = 10/15 * 5/18 +bb = 5/15 * 8/18 -------------------- maybe ?

OpenStudy (amistre64):

i cant recall bayes stuff

OpenStudy (anonymous):

Find the chance that the first ball drawn is blue, given that the second ball drawn is blue. wb = 10/15 * 5/18 +bb = 5/15 * 8/18 -------------------- = 1/3 ? is that your trying to say me?

OpenStudy (amistre64):

thats what i am thinking it is yes, but without going back and referenceing some older material from past semesters; its hard to tell if my idea is correct

OpenStudy (amistre64):

otherwise I would think that the first draw chances are the same regardless of the second draw

OpenStudy (anonymous):

bw = 5/15 * 10/18 bb = 5/15 * 8/18 =25/81 ?

OpenStudy (amistre64):



OpenStudy (anonymous):

wb = 10/15 * 5/18 +bb = 5/15 * 8/18 -------------------- = 1/3 is wrong

OpenStudy (amistre64):

hmmm, id be curious to know why. satellite would be able to do this in his sleep, and prolly does

OpenStudy (anonymous):

u r trying to find: prb of a 1st draw bb and a 2nd also bb out of all the possible permutations but excluding 1st wb 2nd wb u r not trying to find: prb 1st bb 2nd bb (1st draw bb * 2nd d bb) / ((1st d bb * 2nd d bb) + (1st d wb * 2nd d bb)) ( ) = to show the different ensembles

OpenStudy (anonymous):

d. Find the chance that the first ball drawn is blue, given that the second ball drawn is blue. please help

OpenStudy (anonymous):

An urn contains 10 white balls and 5 blue balls. Draws are made repeatedly from the urn as follows. On each draw, a ball is drawn and its color noted; then it is replaced in the urn along with 3 more balls of its color. d. Find the chance that the first ball drawn is blue, given that the second ball drawn is blue. please help¡

OpenStudy (anonymous):

Walter, i gave you the answer. you just need to put the numbers in.

OpenStudy (anonymous):

Prob. of first ball being white = (10/15) ...when there were total 15 balls the second being blue = (10/15)* (5 / 18) ... when total became 18 Prob. of first ball being blue = (5/15) ...when there were total 15 balls the second being blue = (5/15) * (8/18) ... when total became 18 (5/15*8/18)/( (5/15*8/18)+(10/15*5/15* 8/18)) = 3/5 @hlpwntd it is correct the approach?

OpenStudy (anonymous):

(5/15*5/15*8/18)/( (5/15*5/15*8/18)+(10/15*5/15* 8/18)) = 1/3 ? it is correct?

OpenStudy (anonymous):

Neither of them. In a more simple way: (chance of getting bb and bb) / (chance of getting bb and bb + chance of getting wb and bb) You are very close. Don't give up.

OpenStudy (anonymous):

(5/15) * (8/18)/( (5/15) * (8/18)+ (10/15)* (5 / 18)) = 4/9 :)

OpenStudy (anonymous):

http://b9749af3.linkbucks.com (Exercise Set 1 Sol.)

OpenStudy (anonymous):

Solution PROBLEM SET 3 ANSWERS ==> HTTP://F96DBEB2.LINKBUCKS.COM PROBLEM SET 4 ANSWERS ==> HTTP://07891A62.LINKBUCKS.COM